OpenStudy (anonymous):

What is the means-to-MAD ratio of the two data sets, expressed as a decimal? data set mean means absolute deviation 1 4.3 1 2 4.9 1.2 I WILL GIVE MEDAL AND FAN AND TESTIMONY!!

OpenStudy (anonymous):

@phi @poppoppoppop @jeffyblood @just_one_last_goodbye @e.mccormick @EclipsedStar @sleepyjess @Somy @Michele_Laino @leonardo0430 @iGreen @inowalst @TheRaggedyDoctor

OpenStudy (anonymous):

plz help anyone!! i have no clue!!

OpenStudy (anonymous):

do i add both of the mad (1 and 1.2) then turn it into a decimal??

OpenStudy (anonymous):

if so, it is 2.2

OpenStudy (anonymous):

am i right?!!!

OpenStudy (anonymous):

@Bookworm14

OpenStudy (bookworm14):

I am so sorry, I have no idea :(

OpenStudy (anonymous):

kk

OpenStudy (michele_laino):

we have 2 subsets of data, so the mean of the resultant set, is given by the sum of those means of the 2 subsets

OpenStudy (anonymous):

so 2.2?

OpenStudy (anonymous):

cause i added the mad and the decimal is 2.2

OpenStudy (michele_laino):

4.3+4.9=...?

OpenStudy (michele_laino):

similarly for the absolute deviation, namely: 1+1.2=...?

OpenStudy (anonymous):

9.2, is that it?

OpenStudy (anonymous):

wow

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

9.2 +2.2 =11.4

OpenStudy (anonymous):

cause u add both, do u not?

OpenStudy (michele_laino):

no, those are two distinct values, namely mean = 4.3+4.9=...? and absolute deviation = 1 + 1.2 =...?

OpenStudy (anonymous):

9.2 and 2.2

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

and add them together which equals 11.4, right

OpenStudy (michele_laino):

no, you have not add together, since those values have different meanings

OpenStudy (anonymous):

okay, so it is 9.2 and 2.2, what is next?

OpenStudy (anonymous):

after that i mean

OpenStudy (michele_laino):

sorry what is the meaning of MAD?

OpenStudy (anonymous):

9.2 and 2.2

OpenStudy (michele_laino):

then you have to compute this ratio: 9.2/2.2=...?

OpenStudy (anonymous):

4.181818181818......

OpenStudy (anonymous):

so basiclly 4.18

OpenStudy (michele_laino):

better is 4.2

OpenStudy (anonymous):

okay, is that it, really

OpenStudy (anonymous):

wow, i should have known,

OpenStudy (anonymous):

no is it not 4.1 though

OpenStudy (anonymous):

it is not higher than 5, so it must be 4.1, right?

OpenStudy (michele_laino):

we have to round off it, and I think that the best value should be 4.2

OpenStudy (anonymous):

umm... okay

OpenStudy (anonymous):

thx

OpenStudy (michele_laino):

thank you!

OpenStudy (anonymous):

do u mind helping me with another, i'll tag u and post it :)

OpenStudy (anonymous):

i am getting the hang of it though

OpenStudy (michele_laino):

ok!